ELECTRONIC SPECTRA OF BENZENE ISOTOPOMERS IN HELIUM NANODROPLETS

PIERRE CARCABAL, ROMAN SCHMIED, \underline{KEVIN K. LEHMANN}, and GIACINTO SCOLES, Department of Chemistry, Princeton University, Princeton, NJ 08544.

We have recorded high-resolution ultraviolet spectra of four benzene isotopomers (C_{6}H_{6}, C_{6}H_{5}D, 1,3,5-C_{6}H_{3}D_{3}, and C_{6}D_{6}) in helium nanodroplets in the region of the first Herzberg-Teller allowed vibronic transition ^{1}B_{2u} ←^{1}A_{1g} 6_{0}^{1}. The spectra could not be observed by laser-induced fluorescence, but were recorded by beam depletion using bolometric detection. Like tetracene and pentacene,^{a} these benzene isotopomers show multiple absorption lines spread over ∼ 10 cm^{−1}. The structures of these lines are found to be qualitatively different for the various isotopomers. The zero-phonon lines are blue shifted with respect to the gas phase transitions by 3 to 4 cm^{−1}, as predicted by Even et al.^{b}

aM. Hartmann, A. Lindinger, J. P. Toennies, and A. F. Vilesov, Phys. Chem. Chem. Phys. 4, 4839 (2002)

bU. Even, I. Al-Hroub, and J. Jortner, J. Chem. Phys. 115, 2069 (2001)